# Relative equilibria for Scaling Symmetries and Central Configurations

### v1

@@ 7 Relative Equilibria of Scaling Symmetry for Cotangent Bundles
[…]
The Legendre transformation, or fiber derivative in the case of simple mechanical systems is the map $\mathbb{F}L:TQ\to T^{\ast}Q$ defined by the equation
$\left\langle\mathbb{F}L(v_{q}),w_{q}\right\rangle=\left\langle\left\langle v_{q},w_{q}\right\rangle\right\rangle$
for all $q\in Q$ and $v_{q},w_{q}\in T_{q}Q$. In local coordinates we have that $\mathbb{F}L(\mathbf{q},\mathbf{\dot{q}})=\left(\mathbf{q},\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)$.
[…]
$\left\langle\left\langle p_{q},\tilde{p}_{q}\right\rangle\right\rangle_{T^{\ast}Q}=\left\langle\left\langle\mathbb{F}L^{-1}(p_{q}),\mathbb{F}L^{-1}({\tilde{p}}_{q})\right\rangle\right\rangle$
for all $p_{q},\tilde{p}_{q}\in T^{\ast}Q$. The Lagrangian system on $TQ$ corresponds, via the Legendre transform, to a Hamiltonian system on $T^{\ast}Q$ with Hamiltonian of the form $H(p_{q})=K(p_{q})+U(q)$, with
$K(p_{q})=\frac{1}{2}\left\langle\left\langle p_{q},p_{q}\right\rangle\right\rangle_{T^{\ast}Q}=\frac{1}{2}\left\langle\left\langle\mathbb{F}L^{-1}(p_{q}),\mathbb{F}L^{-1}({\tilde{p}}_{q})\right\rangle\right\rangle=K(\mathbb{F}L^{-1}(p_{q})).$
Let $\Psi$ be the action of $G=\mathbb{R}^{+}$ on $Q$, the configuration manifold of a simple mechanical system, and assume that $\Psi$ satisfies the following conditions:
@@
The potential energy $U(q)$ is conformally invariant with parameter $b$, meaning that $\Psi_{g}^{*}U(q)=U(\Psi_{g}(q))=g^{b}U(q)$ for some $b\in\mathbb{R}$.
@@
The kinetic energy $K(v_{q})$ is conformally invariant with parameter $a$, meaning that $\Psi_{g}^{*}K(v_{q})=K((\Psi_{g})_{*}v_{q})=g^{a}K(v_{q})$ for some $a\in\mathbb{R}$.
We have the following Lemma:
Lemma 7.1.
Let $v_{q}\in TQ$ and let $p_{q}=\mathbb{F}L(v_{q})$ be in $T^{\ast}Q$, then
$\mathbb{F}L(T\Psi_{g}(v_{q}))=\hat{\Psi}_{g}(p_{q}),$
where $\hat{\Psi}_{g}$ is the cotangent lift of the diffeomorphism $\Psi_{g}$.
@@
We perform the necessary calculations in local coordinates. Let $(\mathbf{q},\mathbf{v})$ be local coordinates in $TQ$, and let $(\mathbf{\bar{q}},\mathbf{\bar{v}})=T\Psi_{g}(\mathbf{q},\mathbf{v})=(\Psi_{g}(\mathbf{q}),D\Psi_{g}\cdot\mathbf{v})$.
Then, the conjugate momentum transforms as follows:
$\mathbf{\bar{p}}=\frac{\partial L}{\partial\mathbf{v}}\cdot\frac{\partial\mathbf{v}}{\partial\mathbf{\bar{v}}}=\frac{\partial L}{\partial\mathbf{v}}\cdot\frac{\partial\mathbf{v}}{\partial(D\Psi_{g}\cdot\mathbf{v})}=\mathbf{p}\cdot(D\Psi_{g})^{-1},$
if $\mathbf{p}$ is represented as a row vector, or equivalently,
$\mathbf{\bar{p}}=\left((D\Psi_{g})^{-1}\right)^{T}\cdot\mathbf{p},$
if $\mathbf{p}$ is represented as a column vector. Hence $\mathbb{F}L(T\Psi_{g}(\mathbf{q},\mathbf{v}))=(\Psi_{g}(\mathbf{q}),\left((D\Psi_{g})^{-1}\right)^{T}\cdot\mathbf{p})$. This expression is precisely the transformation law for the cotangent lift in local coordinates (see, for instance, [17], page 216). ∎
Proposition 7.2.
Under Assumption 1 and 2, if the action $\Psi$ induces a scaled cotangent lifted action $\hat{\Psi}^{s}$ on $T^{*}Q$, then $\hat{\Psi}^{s}$ is a conformally symplectic action. In particular, if $\hat{\Psi}_{g}^{s}=g^{(a+b)/2}\hat{\Psi}_{g}$, then $\hat{\Psi}_{g}^{s}$ satisfies
$(\hat{\Psi}_{g}^{s})^{*}K(p_{q})=K(\hat{\Psi}_{g}^{s}(p_{q}))=g^{b}K(p_{q}).$
In this case, the Hamiltonian $H$ is scaling invariant, meaning
$(\hat{\Psi}_{g}^{s})^{*}H=g^{b}H.$
Thus, $\hat{\Psi}_{g}$ represents a scaling symmetry.
Proof.
Let $v_{q}\in TQ$ and let $p_{q}=\mathbb{F}L(v_{q})$ be in $T^{\ast}Q$. Under Assumption 2 we have
$\Psi_{g}^{\ast}\left\langle\left\langle v_{q},v_{q}\right\rangle\right\rangle=\left\langle\left\langle T\Psi_{g}(v_{q}),T\Psi_{g}(v_{q})\right\rangle\right\rangle=g^{a}\left\langle\left\langle v_{q},v_{q}\right\rangle\right\rangle.$
On the one hand we have
$g^{a}\left\langle\left\langle v_{q},v_{q}\right\rangle\right\rangle=g^{a}\left\langle\left\langle\mathbb{F}L^{-1}(\mathbb{F}L(v_{q})),\mathbb{F}L^{-1}(\mathbb{F}L(v_{q}))\right\rangle\right\rangle=g^{a}\left\langle\left\langle\mathbb{F}L(v_{q}),\mathbb{F}L(v_{q})\right\rangle\right\rangle_{T^{\ast}Q}=g^{a}\left\langle\left\langle p_{q},p_{q}\right\rangle\right\rangle_{T^{\ast}Q}$
On the other hand by Lemma 7.1 we have
$\left\langle\left\langle T\Psi_{g}(v_{q}),T\Psi_{g}(v_{q})\right\rangle\right\rangle$ $=\left\langle\left\langle\mathbb{F}L^{-1}(\mathbb{F}L(T\Psi_{g}(v_{q}))),\mathbb{F}L^{-1}(\mathbb{F}L(T\Psi_{g}(v_{q})))\right\rangle\right\rangle$
$=\left\langle\left\langle\mathbb{F}L(T\Psi_{g}(v_{q})),\mathbb{F}L(T\Psi_{g}(v_{q}))\right\rangle\right\rangle_{T^{\ast}Q}$
$=\left\langle\left\langle\hat{\Psi}_{g}p_{q},\hat{\Psi}_{g}p_{q}\right\rangle\right\rangle_{T^{\ast}Q}.$
It follows that
$\left\langle\left\langle\hat{\Psi}_{g}p_{q},\hat{\Psi}_{g}p_{q}\right\rangle\right\rangle_{T^{\ast}Q}=g^{a}\left\langle\left\langle p_{q},p_{q}\right\rangle\right\rangle_{T^{\ast}Q}$
Let $\hat{\Psi}_{g}^{s}=g^{s}\hat{\Psi}_{g}$ be the scaled cotangent lift of $\Psi_{g}$, then
$\left\langle\left\langle\hat{\Psi}_{g}^{s}p_{q},\hat{\Psi}_{g}^{s}p_{q}\right\rangle\right\rangle_{T^{\ast}Q}=\left\langle\left\langle g^{s}\hat{\Psi}_{g}p_{q},g^{s}\hat{\Psi}_{g}p_{q}\right\rangle\right\rangle_{T^{\ast}Q}=g^{2s}\left\langle\left\langle\hat{\Psi}_{g}p_{q},\hat{\Psi}_{g}p_{q}\right\rangle\right\rangle_{T^{\ast}Q}=g^{a+2s}\left\langle\left\langle p_{q},p_{q}\right\rangle\right\rangle_{T^{\ast}Q}$
By Assumption 1, we have $\Psi_{g}^{\ast}U=g^{b}U$. If we require $\Psi_{g}^{\ast}H=g^{b}H$, it follows that $a+2s=b$, or equivalently, $s=\frac{a+b}{2}$.
@@
Since the action is conformally symplectic by Proposition 7.2, Theorem 5.4 implies that the momentum map corresponding to the action $\hat{\Psi}^{s}$ is the map $J:T^{*}Q\to\mathfrak{g}^{*}$ given by
$\left\langle J(p_{q}),\eta\right\rangle=\left\langle p_{q},\eta_{Q}(q)\right\rangle,$
where $\eta_{Q}$ is the infinitesimal generator of the action on $Q$ corresponding to $\eta\in\mathfrak{g}$.
For each $q\in Q$ the locked inertia tensor is the map $\mathbb{I}(q):\mathfrak{g}\to\mathfrak{g}^{\ast}$ given by
$\left\langle\eta,\mathbb{I}(q)\xi\right\rangle=\left\langle\left\langle\xi_{Q}(q),\eta_{Q}(q)\right\rangle\right\rangle$
for every $\xi,\eta\in\mathfrak{g}$. Given $\xi\in\mathfrak{g}$ the augmented potential $U_{\xi}:Q\to\mathbb{R}$ is defined by
@@
$K_{\xi}(p_{q})=\frac{1}{2}\|p_{q}-\mathbb{F}L(\xi_{Q}(q))\|_{T^{\ast}Q}^{2}.$
@@
By the definition of $\left\langle\left\langle\cdot,\cdot\right\rangle\right\rangle_{T^{\ast}Q}$ and the definition of locked inertia tensor we have that
$\left\langle\left\langle\mathbb{F}L(\xi_{Q}(q)),\mathbb{F}L(\xi_{Q}(q))\right\rangle\right\rangle_{T^{\ast}Q}=\left\langle\left\langle\xi_{Q}(q),\xi_{Q}q)\right\rangle\right\rangle=\left\langle\xi,\mathbb{I}(q)\xi\right\rangle.$
@@
$\left\langle J(z),\xi\right\rangle$ $=\left\langle p_{q},\xi_{Q}(q)\right\rangle=\left\langle\mathbb{F}L(\mathbb{F}L^{-1}(p_{q})),\xi_{Q}(q)\right\rangle=\left\langle\left\langle\mathbb{F}L^{-1}(p_{q}),\xi_{Q}(q)\right\rangle\right\rangle$
$=\left\langle\left\langle\mathbb{F}L^{-1}(p_{q}),\mathbb{F}L^{-1}(\mathbb{F}L(\xi_{Q}(q)))\right\rangle\right\rangle=\left\langle\left\langle p_{q},\mathbb{F}L(\xi_{Q}(q))\right\rangle\right\rangle_{T^{\ast}Q}$
@@
$K_{\xi}(z)+$ $U_{\xi}(z)=\frac{1}{2}\|p_{q}-\mathbb{F}L(\xi_{Q}(q))\|^{2}_{T^{\ast}Q}+U(q)-\frac{1}{2}\langle\xi,\mathbb{I}(q)\xi\rangle$
$=\frac{1}{2}\|p_{q}\|^{2}_{T^{\ast}Q}-\langle\langle p_{q},\mathbb{F}L(\xi_{Q}(p_{q}))\rangle\rangle_{T^{\ast}Q}+\frac{1}{2}\|\mathbb{F}L(\xi_{Q}(q))\|^{2}_{T^{\ast}Q}+U(q)-\frac{1}{2}\langle\xi,\mathbb{I}(q)\xi\rangle$
$=\frac{1}{2}\|p_{q}\|^{2}_{T^{\ast}Q}-\langle J(p_{q}),\xi\rangle+\frac{1}{2}\langle\xi,\mathbb{I}(q)\xi\rangle+U(q)-\frac{1}{2}\langle\xi,\mathbb{I}(q)\xi\rangle$
$=\frac{1}{2}\|p_{q}\|^{2}_{T^{\ast}Q}-\langle J(p_{q}),\xi\rangle+U(q)$
$=H(p_{q})-\langle J(p_{q}),\xi\rangle$
$=H_{\xi}(z).$
@@
Consider a simple mechanical system on $Q$ with a Lagrangian function $L:TQ\to\mathbb{R}$, and an action of the group $G=\mathbb{R}^{+}$ on $Q$. Assume the conditions stated at the beginning of this section hold. Then, a point $z=p_{q}=(q,p)\in T^{*}Q$ is a relative equilibrium of the scaling symmetry if and only if:
@@
    where $\theta_{L}=(\mathbb{F}L)^{*}\theta_{0}$ is called the Lagrangian one-form (see [21] for more details about the Lagrangian one-form).
@@
Let $(\mathbf{q},\mathbf{p})$ be local coordinates, and let $\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\xi_{Q}}}=\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\dot{q}}=\xi_{Q}(\mathbf{q})}$, then we have that
$K_{\xi}(\mathbf{q},\mathbf{p})=\frac{1}{2}\left\|\mathbf{p}-\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\xi_{Q}}\right\|^{2},\quad H_{\xi}(\mathbf{q},\mathbf{p})=K_{\xi}(\mathbf{q},\mathbf{p})+U_{\xi}(\mathbf{q}),$
and the Liouville one form is $\theta_{0}=\mathbf{p}d\mathbf{q}$. Hence,
@@
$=-\left(\mathbf{p}-\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\xi_{Q}}\right)\left(\frac{\partial^{2}L}{\partial\mathbf{q}\partial\mathbf{\dot{q}}}\right)_{\xi_{Q}}d\mathbf{q}+\left(\mathbf{p}-\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\xi_{Q}}\right)d\mathbf{p}+\frac{\partial U_{\xi}}{\partial\mathbf{q}}d\mathbf{q}+(c\xi)\,\mathbf{p}d\mathbf{q}$
$=\left[-\left(\mathbf{p}-\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\xi_{Q}}\right)\left(\frac{\partial^{2}L}{\partial\mathbf{q}\partial\mathbf{\dot{q}}}\right)_{\xi_{Q}}+\frac{\partial U_{\xi}}{\partial\mathbf{q}}+(c\xi)\,\mathbf{p}\right]d\mathbf{q}+\left(\mathbf{p}-\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\xi_{Q}}\right)d\mathbf{p}=0$
It follows that
$\mathbf{p}=\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\xi_{Q}},\quad\mbox{and}\quad\frac{\partial U_{\xi}}{\partial\mathbf{q}}d\mathbf{q}+(c\xi)\,\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\xi_{Q}}d\mathbf{q}=0.$
Since $\frac{\partial L}{\partial\mathbf{\dot{q}}}d\mathbf{q}$ is the coordinate expression of the Lagrangian one-form $\theta_{L}$ (see [21]), the proof follows.

### v2

@@ 7 Relative Equilibria of Scaling Symmetry for Cotangent Bundles
We now turn our attention to simple mechanical systems, which are defined on a manifold $Q$ with a Lagrangian $L:TQ\to\mathbb{R}$. In this context $(q,v)\in TQ$ is an element of the tangent bundle, $v\in T_{q}Q$ represents a tangent vector in the tangent space at $q$, $(q,p)\in T^{*}Q$ is an element of the cotangent bundle, and $p\in T_{q}^{*}Q$ denotes an element of the cotangent space at $q$.
The Lagrangian is given by:
$L(q,v)=K(q,v)-U(q),$
where $K(q,v)$ represents the kinetic energy, and $U(q)$ denotes the potential energy.
Assume that $Q$ is equipped with a Riemannian metric $g$, which defines a smoothly varying collection of inner products $\langle\langle\cdot,\cdot\rangle\rangle_{q}$ on each tangent space $T_{q}Q$. This inner product induces a norm $\|\cdot\|_{q}$, defined by:
$\|v\|^{2}_{q}=\langle\langle v,v\rangle\rangle_{q},\quad v\in T_{q}Q.$
The kinetic energy is then expressed as:
$K(q,v)=\frac{1}{2}\|v\|_{q}^{2},$
where $v\in T_{q}Q$ is a tangent vector at $q\in Q$.
Let $L:TQ\to\mathbb{R}$ be a smooth Lagrangian function. At each point $q\in Q$, we restrict $L$ to the fiber over $q$, namely the tangent space $T_{q}Q$, giving a function:
$L_{q}:T_{q}Q\to\mathbb{R}.$
On each tangent space, we define the mapping:
$\mathbb{F}L_{q}:T_{q}Q\to T_{q}^{*}Q,\quad v\mapsto dL_{q}(v),$
where $dL_{q}(v)\in T_{q}^{*}Q$ is the derivative (or differential) of $L_{q}$ at $v$. This map associates a velocity $v\in T_{q}Q$ with a momentum $p=dL_{q}(v)\in T_{q}^{*}Q$.
For simple mechanical systems, given $q\in Q$, this fiber-wise map satisfies:
$\left\langle\mathbb{F}L_{q}(v),w\right\rangle=\left.\frac{d}{ds}\right|_{s=0}L(q,v+sw)=\left\langle\left\langle v,w\right\rangle\right\rangle_{q},$
for all $v,w\in T_{q}Q$, where the inner product on the right-hand side is the one induced by the Riemannian metric.
These fiber-wise maps $\mathbb{F}L_{q}$ collectively assemble into the global map known as the Legendre transform or fiber derivative:
$\mathbb{F}L:TQ\to T^{*}Q,\quad(q,v)\mapsto(q,p)=(q,\mathbb{F}L_{q}(v)).$
In local coordinates, the Legendre transformation takes the form:
$\mathbb{F}L(\mathbf{q},\mathbf{\dot{q}})=\left(\mathbf{q},\mathbb{F}L_{q}(\mathbf{q},\mathbf{\dot{q}})\right)=\left(\mathbf{q},\frac{\partial L}{\partial\mathbf{\dot{q}}}\right),$
where $\mathbf{q}$ are the position coordinates and $\mathbf{\dot{q}}$ are the velocity coordinates.
Then the inner product on the fibers of $T^{\ast}Q$ induced by $\left\langle\left\langle\cdot,\cdot\right\rangle\right\rangle_{q}$ is defined by the expression
$\left\langle\left\langle p,\tilde{p}\right\rangle\right\rangle_{q}^{\ast}=\left\langle\left\langle\mathbb{F}L_{q}^{-1}(p),\mathbb{F}L_{q}^{-1}({\tilde{p}})\right\rangle\right\rangle_{q}$
for all $p,\tilde{p}\in T_{q}^{\ast}Q$. The Lagrangian system on $TQ$ corresponds, via the Legendre transform, to a Hamiltonian system on $T^{\ast}Q$ with Hamiltonian:
$H(q,p)=K(q,p)+U(q),$
where:
$K(q,p)=\frac{1}{2}\left\langle\left\langle p,p\right\rangle\right\rangle_{q}^{\ast}=\frac{1}{2}\left\langle\left\langle\mathbb{F}L_{q}^{-1}(p),\mathbb{F}L_{q}^{-1}({p})\right\rangle\right\rangle_{q}.$
Note that we use the same symbol $K$ for both the function $K(q,p)$ defined on the $T^{*}Q$ and $K(q,v)$ defined on $TQ$, which is a slight abuse of notation.
Let $\Psi$ be the action of $\mathbb{R}^{+}$ on $Q$, the configuration manifold of a simple mechanical system, and let $\hat{\Psi}$ be its cotangent lift. Assume that $\Psi$ satisfies the following conditions:
@@
The potential energy $U(q)$ is conformally invariant with parameter $b$ with respect to $\hat{\Psi}$, meaning that $\hat{\Psi}_{g}^{*}U(q)=\Psi_{g}^{*}U(q)=U(\Psi_{g}(q))=g^{b}U(q)$ for some $b\in\mathbb{R}$.
@@
The kinetic energy $K(q,p)$ is conformally invariant with parameter $a$ with respect to $\hat{\Psi}$, meaning that $\hat{\Psi}_{g}^{*}K(q,p)=g^{a}K(q,p)$ for some $a\in\mathbb{R}$.
We have the following proposition:
Proposition 7.1.
Under Assumption 1 and 2, if the action $\Psi$ induces a scaled cotangent lifted action $\hat{\Psi}^{c}$ on $T^{*}Q$, defined by
$\hat{\Psi}^{c}_{g}(q,p)=\left(\Psi_{g}(q),g^{c}\cdot(d\Psi_{g^{-1}})^{\ast}(p)\right),$
then $\hat{\Psi}^{c}$ is a conformally symplectic action. In particular, if $c=(b-a)/2$, then $\hat{\Psi}_{g}^{c}$ satisfies
$(\hat{\Psi}_{g}^{c})^{*}K(q,p)=K(\hat{\Psi}_{g}^{c}(q,p))=g^{b}K(q,p).$
In this case, the Hamiltonian $H$ is scaling invariant, meaning
$(\hat{\Psi}_{g}^{c})^{*}H=g^{b}H.$
Thus, $\hat{\Psi}_{g}^{c}$ represents a scaling symmetry.
@@
The cotangent lifted action $\hat{\Psi}_{g}$ is defined by
$\hat{\Psi}_{g}(q,p)=\left(\Psi_{g}(q),(d\Psi_{g^{-1}})^{\ast}(p)\right).$
Hence, by Assumption 2,
$(\hat{\Psi}_{g})^{\ast}K(q,p)=\frac{1}{2}\left\langle\left\langle(d\Psi_{g^{-1}})^{\ast}(p),(d\Psi_{g^{-1}})^{\ast}(p)\right\rangle\right\rangle_{q}^{\ast}=\frac{1}{2}g^{a}\left\langle\left\langle p,p\right\rangle\right\rangle_{q}^{\ast}.$
On the other hand, if $c=(b-a)/2$, then
$(\hat{\Psi}_{g}^{c})^{\ast}K(q,p)$ $=\frac{1}{2}\left\langle\left\langle g^{c}\cdot(d\Psi_{g^{-1}})^{\ast}(p),g^{c}\cdot(d\Psi_{g^{-1}})^{\ast}(p)\right\rangle\right\rangle_{q}^{\ast}$
$=\frac{1}{2}g^{a+2c}\left\langle\left\langle p,p\right\rangle\right\rangle_{q}^{\ast}$
$=\frac{1}{2}g^{b}\left\langle\left\langle p,p\right\rangle\right\rangle_{q}^{\ast}$
$=g^{b}K(q,p).$
By Assumption 1, we also have that $\hat{\Psi}_{g}^{\ast}U=g^{b}U$. Hence, $(\hat{\Psi}_{g}^{c})^{\ast}H=g^{b}H$. Moreover, it is clear that $\hat{\Psi}^{c}$ is conformally symplectic with parameter $c$, by the definition of scaled cotangent lifted action. ∎
@@
Since the action is conformally symplectic by Proposition 7.1, Theorem 5.4 implies that the momentum map corresponding to the action $\hat{\Psi}^{c}$ is the map $J:T^{*}Q\to\mathbb{R}$ given by
$\left\langle J(q,p),\eta\right\rangle=\left\langle p,\eta_{Q}(q)\right\rangle,$
where $\eta_{Q}$ is the infinitesimal generator of the action on $Q$ corresponding to $\eta\in\mathbb{R}$.
For each $q\in Q$ the locked inertia tensor is the map $\mathbb{I}(q):\mathbb{R}\to\mathbb{R}$ given by
$\left\langle\eta,\mathbb{I}(q)\xi\right\rangle=\left\langle\left\langle\xi_{Q}(q),\eta_{Q}(q)\right\rangle\right\rangle_{q}$
for every $\xi,\eta\in\mathbb{R}$. Given $\xi\in\mathbb{R}$ the augmented potential $U_{\xi}:Q\to\mathbb{R}$ is defined by
@@
$K_{\xi}(q,p)=\frac{1}{2}\left(\|p-\mathbb{F}L_{q}(\xi_{Q}(q))\|_{q}^{\ast}\right)^{2}.$
@@
By the definition of $\left\langle\left\langle\cdot,\cdot\right\rangle\right\rangle_{q}^{\ast}$ and the definition of locked inertia tensor we have that
$\left\langle\left\langle\mathbb{F}L_{q}(\xi_{Q}(q)),\mathbb{F}L_{q}(\xi_{Q}(q))\right\rangle\right\rangle_{q}^{\ast}=\left\langle\left\langle\xi_{Q}(q),\xi_{Q}(q))\right\rangle\right\rangle_{q}=\left\langle\xi,\mathbb{I}(q)\xi\right\rangle.$
@@
$\left\langle J(q,p),\xi\right\rangle$ $=\left\langle p,\xi_{Q}(q)\right\rangle=\left\langle\mathbb{F}L_{q}(\mathbb{F}L_{q}^{-1}(p)),\xi_{Q}(q)\right\rangle=\left\langle\left\langle\mathbb{F}L_{q}^{-1}(p),\xi_{Q}(q)\right\rangle\right\rangle_{q}$
$=\left\langle\left\langle\mathbb{F}L_{q}^{-1}(p),\mathbb{F}L_{q}^{-1}(\mathbb{F}L_{q}(\xi_{Q}(q)))\right\rangle\right\rangle_{q}=\left\langle\left\langle p,\mathbb{F}L_{q}(\xi_{Q}(q))\right\rangle\right\rangle_{q}^{\ast}$
@@
$K_{\xi}(q,p)+$ $U_{\xi}(q,p)=\frac{1}{2}(\|p-\mathbb{F}L_{q}(\xi_{Q}(q))\|^{\ast}_{q})^{2}+U(q)-\frac{1}{2}\langle\xi,\mathbb{I}(q)\xi\rangle$
$=\frac{1}{2}(\|p\|^{\ast}_{q})^{2}-\langle\langle p,\mathbb{F}L_{q}(\xi_{Q}(q))\rangle\rangle_{q}^{\ast}+\frac{1}{2}(\|\mathbb{F}L_{q}(\xi_{Q}(q))\|^{\ast}_{q})^{2}+U(q)-\frac{1}{2}\langle\xi,\mathbb{I}(q)\xi\rangle$
$=\frac{1}{2}(\|p\|^{\ast}_{q})^{2}-\langle J(q,p),\xi\rangle+\frac{1}{2}\langle\xi,\mathbb{I}(q)\xi\rangle+U(q)-\frac{1}{2}\langle\xi,\mathbb{I}(q)\xi\rangle$
$=\frac{1}{2}(\|p\|^{\ast}_{q})^{2}-\langle J(q,p),\xi\rangle+U(q)$
$=H(q,p)-\langle J(q,p),\xi\rangle$
$=H_{\xi}(q,p).$
@@
Consider a simple mechanical system on $Q$ with a Lagrangian function $L:TQ\to\mathbb{R}$, and let $\Psi$ be an action of the multiplicative group $\mathbb{R}^{+}$ on $Q$ satisfying Assumptions 1 and 2. This action induces a scaled cotangent lift $\hat{\Psi}^{c}$ with $c=(b-a)/2$, which is a scaling symmetry.
Then, a point $z=(q,p)\in T^{*}Q$ is a relative equilibrium of the scaling symmetry if and only if the following conditions are satisfied:
@@
    where $\theta_{L}=(\mathbb{F}L)^{*}\theta_{0}$ is known as the Lagrangian one-form. For further details on the Lagrangian one-form, see [20].
@@
First notice that $\hat{\Psi}^{c}$ is a scaling symmetry by Proposition 7.1. Then, let $(\mathbf{q},\mathbf{p})$ be local coordinates, and define
$\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\xi}_{Q}}=\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\dot{q}}=\mathbf{\xi}_{Q}(\mathbf{q})},$
where $\mathbf{\xi}_{Q}$ is the infinitesimal generator of the group action. Then we have
$K_{\xi}(\mathbf{q},\mathbf{p})=\frac{1}{2}\left\|\mathbf{p}-\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\xi}_{Q}}\right\|^{2}_{q},$
where $\|\cdot\|_{q}$ is the norm induced by the Riemannian metric, which depends on $\mathbf{q}$. The Hamiltonian is then expressed as
$H_{\xi}(\mathbf{q},\mathbf{p})=K_{\xi}(\mathbf{q},\mathbf{p})+U_{\xi}(\mathbf{q}),$
where $U_{\xi}(\mathbf{q})$ is the amended potential. Furthermore, the Liouville one-form is $\theta_{0}=\mathbf{p}\,d\mathbf{q}$.
Since the following formula holds (see Theorem 3.6 in [12]):
$\frac{\partial}{\partial q_{k}}\left\langle\left\langle u,v\right\rangle\right\rangle_{q}=\left\langle\left\langle\nabla_{\partial_{q_{k}}}u,v\right\rangle\right\rangle_{q}+\left\langle\left\langle u,\nabla_{\partial_{q_{k}}}v\right\rangle\right\rangle_{q},$
where $\nabla$ is the Levi-Civita covariant derivative, it follows that the derivative of $K_{\xi}(\mathbf{q},\mathbf{p})$ with respect to $q_{k}$ is:
$\frac{\partial}{\partial q_{k}}K_{\xi}(\mathbf{q},\mathbf{p})=\frac{1}{2}\frac{\partial}{\partial q_{k}}\left\|\mathbf{p}-\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\xi}_{Q}}\right\|^{2}_{q}=\left\langle\left\langle\nabla_{\partial{q_{k}}}\left(\mathbf{p}-\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\xi}_{Q}}\right),\mathbf{p}-\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\xi}_{Q}}\right\rangle\right\rangle_{q}.$ (13)
Moreover,
$\frac{\partial}{\partial p_{k}}K_{\xi}(\mathbf{q},\mathbf{p})$ $=\frac{1}{2}\frac{\partial}{\partial p_{k}}\left\|\mathbf{p}-\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\xi}_{Q}}\right\|^{2}_{q}$ (14)
$=\left\langle\left\langle\frac{\partial}{\partial p_{k}}\left(\mathbf{p}-\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\xi}_{Q}}\right),\mathbf{p}-\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\xi}_{Q}}\right\rangle\right\rangle_{q}$
$=\left\langle\left\langle\frac{\partial}{\partial p_{k}}\mathbf{p},\mathbf{p}-\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\xi}_{Q}}\right\rangle\right\rangle_{q}$
To compute the differential of the Hamiltonian, we have:
@@
$=\left[\frac{\partial K_{\xi}}{\partial\mathbf{q}}+\frac{\partial U_{\xi}}{\partial\mathbf{q}}+(c\xi)\,\mathbf{p}\right]d\mathbf{q}+\frac{\partial K_{\xi}}{\partial\mathbf{p}}\,d\mathbf{p}$
From which we obtain the equations
$\frac{\partial K_{\xi}}{\partial\mathbf{p}}=0,\quad\frac{\partial K_{\xi}}{\partial\mathbf{q}}+\frac{\partial U_{\xi}}{\partial\mathbf{q}}+(c\xi)\,\mathbf{p}=0$
By equation (14), $\frac{\partial K_{\xi}}{\partial\mathbf{p}}=0$ implies $\mathbf{p}-\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\xi}_{Q}}=0$. Using equation (13) this implies that $\frac{\partial K_{\xi}}{\partial\mathbf{q}}=0$.
From this, it follows that
$\mathbf{p}=\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\xi}_{Q}},\quad\text{and}\quad\frac{\partial U_{\xi}}{\partial\mathbf{q}}+(c\xi)\,\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\xi}_{Q}}=0.$
Multiplying by $d\mathbf{q}$ the second equation can also be written as $\frac{\partial U_{\xi}}{\partial\mathbf{q}}d\mathbf{q}+(c\xi)\,\left(\frac{\partial L}{\partial\mathbf{\dot{q}}}\right)_{\mathbf{\xi}_{Q}}d\mathbf{q}=0$. Since $\frac{\partial L}{\partial\mathbf{\dot{q}}}\,d\mathbf{q}$ is the coordinate expression of the Lagrangian one-form $\theta_{L}$ (see [20]), the proof follows. ∎